# Exotic PeVatrons as sources of ultra-high-energy gamma rays

### v1

@@ I Introduction
The most efficient particle acceleration occurs in uniform electric fields $\mathcal{E}$, where the energy gain as
@@ II Black Holes, Vortices and Magnetic Flux Structures
We begin with the Abelian Higgs model in curved spacetime with action given by
$S=S_{EH}+\int d^{4}x\sqrt{-g}\left[\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+D_{\mu}\Phi^{\dagger}D^{\mu}\Phi-V(\Phi^{\dagger},\Phi)\right],$ (10)
where $S_{EH}$ is the Einstein-Hilbert action, $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the $U(1)$ field strength and $D_{\mu}=\nabla_{\mu}+iqeA_{\mu}$ is the covariant derivative with the electric charge $qe$. The scalar field can be decomposed as
$\Phi(x)=vX(x)e^{i\chi(x)},\quad A_{\mu}(x)=\frac{1}{qe}[P_{\mu}(x)-\nabla_{\mu}\chi(x)]\,.$ (11)
This system admits a Nielsen-Olsen vortex solution. This can be find solving the Equations of Motions in cylindrical coordinates ($\Phi=vX_{0}(R)e^{i\phi}$) become:
@@
where $\beta=\lambda/2e^{2}=m_{scalar}^{2}/m_{vector}^{2}$. While general solutions of these non-linear equations are complicated, their asymptotic behaviors clearly show that vortex solutions with a non-zero topological winding number interpolate two different vacua. In the limit of $R\rightarrow\infty$, a Nielsen-Olsen vortex solution is obtained.
@@
  • For $q\ll 1$ ($M\gg 1$), large magnetic fluxes are possible.
@@ III.1 Boson Stars
where $R$ and $M$ denote the star’s radius and mass, respectively [^1] While Chavanis:2011zi identifies an additional branch for attractive interactions ($\lambda<0$), we exclude this unstable solution from our rotating case analysis..
@@
The winding number dependence follows from Davidson:2016uok , where $\alpha=3/2$ and $\beta=3$ for large $n$. Numerical studies Ontanon:2021hbg confirm the proportionality $M_{\text{crit}}\propto n$, supporting our choice of $\alpha=\beta=2$ for the linear approximation. This simplification agrees with earlier treatments Lee:1986ts ; Ferrell:1989kz and suffices for qualitative analysis of millicharged dark matter BSs.

### v2

@@ I Introduction
The most efficient particle acceleration occurs in uniform electric fields $\mathcal{E}$, where the energy of a nucleus with electric charge $Ze$ gains as
@@ II Black Holes, Vortices and Magnetic Flux Structures
We begin with the Abelian Higgs model in curved spacetime with action given by Achucarro:1995nu
$S=S_{EH}+\int d^{4}x\sqrt{-g}\left[\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+D_{\mu}\varphi^{\dagger}D^{\mu}\varphi-V(\varphi^{\dagger},\varphi)\right],$ (10)
where $S_{EH}$ is the Einstein-Hilbert action, $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the $U(1)$ field strength, $D_{\mu}=\nabla_{\mu}+iqeA_{\mu}$ is the covariant derivative with the electric charge $qe$ and $\nabla_{\mu}$ is the covariant derivative with respect to the metric $g_{\mu\nu}$ in curved spacetime. The scalar field can be decomposed as
$\varphi(x)=vX(x)e^{i\chi(x)},\quad A_{\mu}(x)=\frac{1}{qe}[P_{\mu}(x)-\nabla_{\mu}\chi(x)]\,.$ (11)
This system admits a Nielsen-Olsen vortex solution. This can be find solving the Equations of Motions in cylindrical coordinates ($\varphi(x)=vX_{0}(R)e^{i\phi}$) become:
@@
where $\beta=\lambda/2e^{2}=m_{scalar}^{2}/m_{vector}^{2}$. While general solutions of these non-linear equations are complicated, their asymptotic behaviors clearly show that vortex solutions with a non-zero topological winding number interpolate two different vacua. In the limit of $R\rightarrow\infty$, a Nielsen-Olsen vortex solution is obtained. Note that $\chi$ given in Eq. 11 is a gauge degree of freedom, which is not dynamical but rather a topological quantity characterizing the winding number of the vortex, i.e. $\oint d\chi=2\pi n,(n\in\mathbb{Z})$. Eqs. (12-13) are written in rescaled coordinates with replacement $P_{\mu}\to nP_{\mu}$, representing that the string has width of order unity and winding number one (see ref. Achucarro:1995nu for more details).
@@
  • For $q\ll 1$ ($M\gg 1$), large magnetic fluxes are possible. In literature, dark matter candidates with a tiny fractional electric charge $q\ll 1$ have been proposed, known as millicharged dark matter. Here $q$ is dimensionless and defined with respect to an electric unity charge $e$.
@@ III.1 Boson Stars
where $R$ and $M$ denote the star’s radius and mass, respectively. Inspired by the mass-radius relation of rotating Axion star given in ref. Davidson:2016uok , the parameters $\alpha$ and $\beta$ are introduced by hand in Eq. 25 to account for the rotating states of the BSs. When $n=0$, Eq. 25 reduces to the non-rotating case of the BSs’ mass-radius relation obtained in ref. Chavanis:2011zi . The appropriate values of $\alpha$ and $\beta$ will be determined in large n limit, as discussed later. In addition, we take $\lambda\geqslant 0$. While ref. Chavanis:2011zi identifies an additional branch for attractive interactions ($\lambda<0$), we exclude this unstable solution from our rotating case analysis.
@@
The winding number dependence follows from Davidson:2016uok , where $\alpha=3/2$ and $\beta=3$ for large $n$. Numerical studies Ontanon:2021hbg confirm the proportionality $M_{\text{crit}}\propto n$, supporting another choice of $\alpha=\beta=2$ for the linear approximation. This simplification agrees with earlier treatments Lee:1986ts ; Ferrell:1989kz . Since we perform qualitative analysis of millicharged dark matter BSs, thus we simply take $\alpha=\beta=2$ in the numerical calculation.